# Federated Learning with Reduced Information Leakage and Computation

### v1

@@ 6 Experiments
Convergence and Heterogeneity. We evaluate the convergence rate, training time, and accuracy of Upcycled-FL under different parameter settings. We take $\lambda_{m}=\lambda,\forall m$ in experiments, while we observe similar results for time-varying $\lambda_{m}$. Table 1 compares the training/testing accuracy of different algorithms when the device heterogeneity is high (90% stragglers). The results show that Upcycled-FL outperforms both FedProx and FedAvg on all datasets (with appropriate $\mu$). While FedAvg achieves good performance on Synthetic(iid), it is not robust to heterogeneous data. When data is i.i.d., adding the proximal term $\frac{\mu}{2}||\omega-\overline{\omega}^{t}||^{2}$ may hurt the performance. However, the proximal term can help stabilize the algorithm and significantly improve the performance in practical settings when data is heterogeneous; these observations are consistent with [2]. Importantly, Upcycled-FL is more robust to the level of statistical heterogeneity than FedProx as it could attain consistent improvements for all settings. In practice, $\mu$ needs to be carefully tuned; with a carefully selected $\mu$, Upcycled-FL could achieve $2\%\sim 7\%$ improvement on accuracy compared to FedProx.
Figure 2(a) and 2(b) illustrate the convergence property of Upcycled-FL on Synthetic(0.5,0.5) and Femnist (in comparison with FedProx and FedAvg), and examine the impact of hyper-parameter $\lambda_{m}$. Loss in $y$-axes indicates the averaged loss over all devices. In each iteration, 30% of devices are selected with 90% stragglers. We observe that Upcycled-FL could attain a comparable performance as FedProx in the first 80 (resp. 240) iterations on Synthetic(0.5,0.5) (resp. Femnist), with 50% (resp. 20%) of SGD computation reduced. Specifically, the final training (testing) accuracy attained on Synthetic(0.5,0.5) is 82.61% (82.64%) for Upcycled-FL, 79.55% (79.46%) for FedProx, and 79.53% (78.97%) for FedAvg; while the accuracy attained on Femnist is 81.94% (82.12%) for Upcycled-FL, 76.10% (76.66%) for FedProx, and 20.43% (20.01%) for FedAvg. More results on other datasets are presented in Appendix H.4.
(a) Synthetic(iid)
(b) Synthetic(0,0)
(c) Synthetic(iid)
(d) Synthetic(0,0)
Figure 3: Comparison of private Upcycled-FL and private FedProx using objective perturbation ((a)(b)) and output perturbation ((c)(d)) with $\delta=0.001$. We take $\mu=0.1,\lambda=0.04$ for all datasets. $\bar{\epsilon}$ denotes the average $\epsilon$ among all devices. In both experiments, the privacy of private Upcycled-FL is strictly stronger than private FedProx.
The effect of hyper-parameter $\lambda_{m}=\lambda$ on the convergence rate of Upcycled-FL is also evaluated in Figure 2(a)2(b) (right). Although the sufficient conditions in Theorem 4.6 and Corollary 4.8 suggest that Upcycled-FL converges when $\lambda$ is sufficiently large, the experiments indicate that Upcycled-FL can easily converge in practice even with small $\lambda$. Furthermore, we observe that models with smaller $\lambda$ tend to converge faster. This is because that $\lambda$ controls step size $\frac{\mu}{\mu+\lambda}$ (by Eqn. (5)): a larger $\lambda$ results in a smaller step size and thus slower convergence.
@@
Privacy-Accuracy Tradeoff. We next inspect the accuracy-privacy trade-off of private Upcycled-FL and compare it with private FedProx. We adopt both objective and output perturbation to preserve privacy in the experiments, while other privacy-preserving techniques can also be used. For each parameter setting, we perform 10 independent runs of the experiment and record both the mean and range of their losses. To precisely quantify privacy, we focus on settings without device heterogeneity (no straggler) and low statistical heterogeneity (Synthetic(iid), Synthetic(0,0)).
Figure 3(a) and 3(b) demonstrate the performance (averaged loss and the respective privacy loss $\epsilon$) of (non-)private Upcylced-FL and private FedProx on synthetic data using objective perturbation. As expected, private Upcycled-FL is much more stable and significantly improves both privacy and accuracy over private FedProx. This is because the private Upcycled-FL with less information leakage requires much less perturbation to attain the same privacy guarantee as private FedProx. In general, we observe that private Upcycled-FL can experience 9.5% less privacy loss than private FedProx, and meanwhile having 11.0% (or 11.5%) higher training (or testing) accuracy on average and reducing 48% of training time. We also conduct similar experiments using output perturbation (Figure 3(c) and 3(d)). Results show that on average, private Upcycled-FL can experience 12.0% less privacy loss than private FedProx while having 6.1% (or 6.2%) higher training (or testing) accuracy; the training time is still reduced by 48% under Upcycled-FL.
We also conduct experiments to examine the accuracy-privacy trade-off under different magnitudes of perturbation, the results are presented in Appendix H.5.1. It shows that private Upcycled-FL can consistently outperform private FedProx under different levels of privacy.
@@ 7 Conclusion
This paper proposes Upcycled-FL, a novel federated learning framework under which information leakage and computation cost can be reduced significantly. We theoretically examined the convergence (rate) of Upcycled-FL. The experiments show that compared to existing algorithms, Upcycled-FL can be more robust on heterogeneous data and attain much better privacy-accuracy trade-off under common differential privacy mechanisms.
@@ Appendix A Notation Table
$\mu$ hyper-parameter for proximal term in FedProx and Upcycled-FL
@@
Appendix B Related Work
Differentially private federated learning. Differential privacy has been widely used in federated learning to provide privacy guarantees [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20]. For example, [7] uses the Gaussian mechanism for a federated learning problem and propose an incentive mechanism to encourage users to share their data and participate in the training process. [8] introduces $f$-differential privacy, a generalized version of Gaussian differential privacy, and propose a federated learning algorithm satisfying this new notion. [9] proposes a new mechanism called Random Response with Priori (RRP) to achieve local differential privacy and apply this mechanism to the text data by training a Latent Dirichlet Allocation (LDA) model using a federated learning algorithm. [20] adapts the Bayesian privacy accounting method to the federated setting and propose joint accounting method for estimating client-level and instance-level privacy simultaneously and securely. [19] presents a private scheme that adds noise to parameters at the random selected devices before aggregating and provides a convergence bound. [10] combines the Gaussian mechanism with gradient clipping in federated learning to improve the privacy-accuracy tradeoff. [5] considers a different setting where only the last update is publicly released and the central server and other devices are assumed to be trustworthy. However, all these algorithms need client’s local data to update the model and the information leakage inevitably occurs at every iteration. This is fundamentally different from Upcycled-FL, which reuses existing results to update half of iterations and significantly reduces information leakage and computation.
Tackling heterogeneity in federated learning. It’s worth mentioning that, empirically Upcycled-FL also outperforms existing baseline algorithm under device and statistical heterogeneity. In real-world scenarios, local data are often non-identically distributed across different devices; different devices are also often equipped with different specifications and computation capabilities. Such heterogeneity often causes instability in the model performance and leads to divergence. Many approaches have been proposed to tackle this issue. For example, FedAvg [44] uses a random selection of devices at each iteration to reduce the negative impact of statistical heterogeneity; however, it may fail to converge when heterogeneity increases. Other methods include FedProx [2], a generalization and re-parameterization of FedAvg that adds a proximal term to the objective function to penalize deviations in the local model from the previous aggregation, and FedNova [3] that re-normalizes local updates before updating to eliminate objective inconsistency. It turns out that Upcycled-FL exhibits superior performance in the presence of heterogeneity because gradients encapsulate information on data heterogeneity, the reusing of which leads to a boost in performance.
Appendix C Privacy analysis and results
We provide the privacy analysis of private Upcycled-FL under output and objective perturbation.
Here, we focus on settings where local loss function $F_{i}(\omega_{i},\mathcal{D}_{i}):=\frac{1}{|\mathcal{D}_{i}|}\sum_{d\in\mathcal{D}_{i}}\hat{F}_{i}(\omega_{i},d)$ for some $\hat{F}_{i}$, and the guarantee of privacy is presented in Theorem C.1 (output perturbation) and C.2 (objective perturbation) below. The total privacy loss in the following theorem is composed using moments accountant method [27].
Theorem C.1.
Theorem C.2.
Consider the private Upcycled-FL over $2M$ iterations under objective perturbation with noise $n^{m}_{i}\sim\exp{(-\alpha_{i}^{m}||n_{i}^{m}||_{2})}$. Suppose $\hat{F}_{i}$ is generalized linear model [48, 39]In supervised learning, the sample $d=(x,y)$ corresponds to the feature and label pair. Function $\hat{F}_{i}(\omega,d)$ is generalized linear model if it can be written as a function of $\omega^{T}x$ and $y$. that satisfies $||\nabla\hat{F}_{i}(\omega;d)||<u_{1},~{}|\hat{F}^{\prime\prime}_{i}|\leq u_{2}$. Let feature vectors be normalized such that its norm is no greater than 1, and suppose $u_{2}\leq 0.5|\mathcal{D}_{i}|\mu$ holds. Then the algorithm satisfies $(\varepsilon,0)$-DP for agent $i$ where $\varepsilon=\sum_{m=0}^{M}\frac{2\alpha^{m}_{i}u_{1}\mu+2.8u_{2}}{|\mathcal{D}_{i}|\mu}.$
The assumptions on $\hat{F}_{i}$ are again fairly standard in the literature, see e.g., [36, 42, 43]. Theorem C.1 and C.2 show that the total privacy loss experienced by each agent accumulates over iterations and privacy loss only comes from odd iterations. In contrast, if consider differentially private FedProx, accumulated privacy loss would come from all iterations. Therefore, to achieve the same privacy guarantee, private Upcycled-FL requires much less perturbation per iteration than private FedProx. As a result, accuracy can be improved significantly. Experiments in Section 6 show that Upcycled-FL significantly improves privacy-accuracy trade-off compared to other methods.
Appendix D Lemmas
Lemma D.1.
Define $\widetilde{\omega}^{t}=\mathbb{E}_{i}(\omega_{i}^{t})$. Suppose conditions in Theorem 4.6 hold. Then the following holds for all $m$:
@@ Appendix D Lemmas
Lemma D.2.
Let $\mathcal{S}_{m}$ be the set of $K$ randomly selected local devices got updated at iterations $12m-$ 1 and $2m$, and $\mathbb{E}_{\mathcal{S}_{m}}[\cdot]$ be expectation with respect to the choice of devices. Then we have
@@
Appendix E Proofs
Proof of Lemma D.1
@@ Appendix E Proofs
The Lemma D.1 is proved. ∎
Proof of Lemma D.2
@@
$\mathbb{E}_{\mathcal{S}_{m}}[f(\overline{\omega}^{2m+1})]$
$\leq f(\widetilde{\omega}^{2m+1})$
@@
$\mathbb{E}_{\mathcal{S}_{m}}[f(\overline{\omega}^{2m+1})]$
$\leq f(\widetilde{\omega}^{2m+1})+\frac{2}{K}\frac{L^{2}}{\rho^{2}}||\overline{\omega}^{2m}-\overline{\omega}^{2m-1}||^{2}+\frac{2B^{2}}{K\rho^{2}}||\nabla f(\overline{\omega}^{2m-1})||^{2}$
@@
Lemma 4.5 can be proved by combing Lemmas D.1 and D.2, where
@@
Appendix F Proof of Theorem 4.6
@@
Appendix G Proof of Theorem C.1
@@
According to [27], for a mechanism $\mathscr{M}$ outputs $o$, with inputs $d$ and $\hat{d}$, let a random variable $c(o;\mathscr{M},d,\hat{d})=\log\frac{\text{Pr}(\mathscr{M}(d)=o)}{\text{Pr}(\mathscr{M}(\hat{d})=o)}$ denote the privacy loss at $o$, and
@@
Use the tail bound [Theorem 2, [27]], for any $\varepsilon\geq\frac{M\tau^{2}}{2|D|^{2}\sigma^{2}}$, the algorithm is $(\varepsilon,\delta)$-differentially private for
@@
To find $\lambda^{*}=\underset{\lambda:\lambda\geq 0}{\text{argmin}}~{}h(\lambda)$, take derivative of $h(\lambda)$ and assign 0 gives the solution $\bar{\lambda}=\frac{\varepsilon|D|^{2}\sigma^{2}}{M\tau^{2}}-\frac{1}{2}\geq 0$, and $h^{\prime\prime}(\bar{\lambda})>0$, implies $\lambda^{*}=\bar{\lambda}$. Plug into (G) gives:
@@
Theorem C.1 is proved. ∎
Appendix H Experiments
H.1 Details of Datasets
@@
Synthetic. The synthetic data is generated using the same method in [2]. We briefly describe the generating steps here. For each device k, $y_{k}$ is computed from a softmax function $y_{k}=\text{argmax}(\text{softmax}(W_{k}x_{k}+b_{k}))$. $W_{k}$ and $b_{k}$ are drawn from the same Gaussian distribution with mean $u_{k}$ and variance $1$, where $u_{k}\in N(0;\beta)$. $x_{k}\in N(v_{k};\Sigma)$. $v_{k}$ is drawn from a Gaussian distribution with mean $B_{k}\in\mathcal{N}(0,\gamma)$ and variance $1$. $\Sigma$ is diagonal with $\sum j,j=j^{-1.2}$. In such a setting, $\beta$ controls how many local models differ from each other and $\gamma$ controls how much local data at each device differs from that of other devices.
In our experiment, we take $k=30$, $x\in\mathcal{R}^{20}$, $W\in\mathcal{R}^{10*20}$, $b\in\mathcal{R}^{10}$. We generate 4 datasets in total. They’re Synthetic(iid) Synthetic(0,0) with $\beta=0$ and $\gamma=0$, Synthetic(0.5,0.5) with $\beta=0.5$ and $\gamma=0.5$ and Synthetic(1,1) with $\beta=1$ and $\gamma=1$.
Femnist: Similar with [2], we subsample 10 lower case characters (‘a’-‘j’) from EMNIST [45] and distribute 5 classes to each device. There are 50 devices in total. The input is 28*28 image.
Sent140: A text sentiment analysis task on tweets [46]. The input is sequence of length 25 and output is the probabilities of 2 classes.

### v2

@@ 7.3 Results
Because even iterations of Upcycled-FL only involve addition/subtraction operations with no transmission overhead and almost no computational cost, we train the Upcycled version of FL algorithms with double iterations compared to baselines in approximately same training time in this experiment. We evaluate the convergence rate and accuracy of Upcycled-FL under different dataset and heterogeneity settings. In each iteration, 30% of devices are selected with 90% stragglers. Table 1 compares the average accuracy of different algorithms when the device heterogeneity is high (90% stragglers). The results show that all baselines can be enhanced by our methods. Notably, while FedAvg achieves good performance on Syn(iid), it is not robust on heterogeneous data, e.g. Syn(0,0), Syn(0.5,0.5), and Syn(1,1). Nonetheless, Upcycled-FedAvg makes it comparable with the regular FedProx algorithm, which shows that our strategy can also mitigate performance deterioration induced by data heterogeneity. When data is i.i.d., FedProx with the proximal term $\frac{\mu}{2}||\omega-\overline{\omega}^{t}||^{2}$ may hurt the performance compared with FedAvg. However, the proximal term can help stabilize the algorithm and significantly improve the performance in practical settings when data is heterogeneous; these observations are consistent with (Li et al., 2020). Importantly, Upcycled-FL strategy further makes it more robust to statistical heterogeneity than regular FedProx as it could attain consistent improvements for all settings.
(a) Syn(iid)
(b) Syn(0.5,0.5)
(c) FEMNIST
Figure 2: Comparison on average loss and standard deviation between Upcycled-FL methods and original FL algorithms in the non-private setting under the approximate same training time. The training time refers to the time needed for a given number of iterations. Upcycled-FL does not require an update in the even iterations, allowing Upcycled-FL to train with doubled iterations.
(a) Syn(iid)
(b) Syn(0,0)
(c) Syn(0.5,0.5)
(d) Syn(1,1)
Figure 3: Comparison on average loss and standard deviation of private Upcycled-FL and private FL methods using output perturbation. The noise parameter $\sigma$ is 1.0 for all baselines, while $\sigma$ of the Upcycled version is set to 0.8. Taking the iid dataset as an example, $\bar{\epsilon}=1.40$ for the Upcycled version, which ensures stronger privacy than the original method with $\bar{\epsilon}=1.59$.
Figure 2(a), 2(b) and 2(c) compare the convergence property on Syn(iid), Syn(0.5,0.5) and FEMNIST. Note that, by using the aggregation rule in Figure 1(b), the training time of Upcycled-FL is almost the same as the baselines without introducing extra cost. We observe that under the same training time (the number of iterations for baselines), Upcycled-FL strategy brings benefits (achieving lower loss along training) for baselines in most cases. And this improvement is significant when the dataset is i.i.d. The loss trends are consistent with results in Table 1. We provide more results on the other three datasets in Appendix F.3.
(a) Syn(iid)
(b) Syn(0,0)
(c) Syn(0.5,0.5)
(d) Syn(1,1)
Figure 4: Comparison on average loss and standard deviation of private Upcycled-FL and private FL methods using objective perturbation. Under objective perturbation, noise parameter $\alpha$ is 10 for all baselines, while $\alpha$ of the Upcycled version is set to 20 to ensure stronger privacy than the original versions. Taking the iid dataset as an example, $\bar{\epsilon}$ associated with these noise parameters is 7.36 for FedProx and 7.25 for Upcycled-FedProx (when $\mu$ = 0.5).
Privacy-Accuracy Trade-off.
@@
We next inspect the accuracy-privacy trade-off of private Upcycled-FL and compare it with private baselines. Although we adopt both objective and output perturbation to achieve differential privacy, other techniques can also be used. For each parameter setting, we still conduct a grid search and perform 4 independent runs of experiments. To precisely quantify privacy, no straggler is considered in this experiment. We reported the results using output perturbation and objective perturbation. Figure 3 and 4 demonstrate the performance of private Upcylced-FL and private baselines on synthetic data using two types of perturbation respectively. We also report the results on real data in Appendix F.4.
Here we carefully set the perturbation strength of each algorithm such that the privacy loss $\epsilon$ for private Upcycled-FL is strictly less than the original methods. As expected, private Upcycled-FL is more stable than baselines and attains a lower loss value under smaller $\epsilon$. This is because the private Upcycled-FL with less information leakage requires much less perturbation to attain the same privacy guarantee as FL methods under privacy constraints. We also observe that in general Upcycled-FL can be used to augment all baseline methods with or without client local heterogeneity.
@@ 8 Conclusion
This paper proposes Upcycled-FL, a novel plug-in federated learning strategy under which information leakage and computation costs can be reduced significantly. We theoretically examined the convergence (rate) of Upcycled-FedProx, a special case when Upcycled-FL strategy is applied to a well-known FL algorithm named FedProx. Extensive experiments on synthetic and read data further show that Upcycled-FL can be combined with common FL algorithms and enhance their robustness on heterogeneous data while attaining much better privacy-accuracy trade-off under common differential privacy mechanisms.
Acknowledgments
This material is based upon work supported by the U.S. National Science Foundation under awards IIS-2040800, IIS-2112471, IIS-2202699, IIS-2301599, and CMMI-2301601, by grants from the Ohio State University’s Translational Data Analytics Institute and College of Engineering Strategic Research Initiative.
@@ Appendix A Notation Table
$\mu$ hyper-parameter for proximal term in FedProx and Upcycled FedProx
@@
$\tau$ the clipping threshold for output perturbation
Appendix B Lemmas
Lemma B.1.
Define $\widetilde{\omega}^{t}=\mathbb{E}_{i}(\omega_{i}^{t})$. Suppose conditions in Theorem 5.6 hold, then the following holds $\forall m$:
@@ Appendix B Lemmas
Lemma B.2.
Let $\mathcal{S}_{m}$ be the set of $K$ randomly selected local devices got updated at iterations $12m-$ 1 and $2m$, and $\mathbb{E}_{\mathcal{S}_{m}}[\cdot]$ be the expectation with respect to the choice of devices. Then we have
@@
Appendix C Proofs
Proof of Lemma B.1
@@ Appendix C Proofs
The Lemma B.1 is proved. ∎
Proof of Lemma B.2
@@
$\mathbb{E}_{\mathcal{S}_{m}}[f(\overline{\omega}^{2m+1})]\leq$ $f(\widetilde{\omega}^{2m+1})$
@@
$\mathbb{E}_{\mathcal{S}_{m}}[f(\overline{\omega}^{2m+1})]\leq$ $f(\widetilde{\omega}^{2m+1})+\frac{2}{K}\frac{L^{2}}{\rho^{2}}||\overline{\omega}^{2m}-\overline{\omega}^{2m-1}||^{2}+\frac{2B^{2}}{K\rho^{2}}||\nabla f(\overline{\omega}^{2m-1})||^{2}$
@@
Lemma 5.5 can be proved by combing Lemmas B.1 and B.2, where
@@
Appendix D Proof of Theorem 5.6
@@
Appendix E Proof of Theorem 6.2
@@
According to Abadi et al. [2016], for a mechanism $\mathscr{M}$ outputs $o$, with inputs $d$ and $\hat{d}$, let a random variable $c(o;\mathscr{M},d,\hat{d})=\log\frac{\text{Pr}(\mathscr{M}(d)=o)}{\text{Pr}(\mathscr{M}(\hat{d})=o)}$ denote the privacy loss at $o$, and
@@
Use the tail bound [Theorem 2, Abadi et al. [2016]] , for any $\varepsilon\geq\frac{M\tau^{2}}{2|D|^{2}\sigma^{2}}$, the algorithm is $(\varepsilon,\delta)$-differentially private for
@@
To find $\lambda^{*}=\underset{\lambda:\lambda\geq 0}{\text{argmin}}~{}h(\lambda)$, take derivative of $h(\lambda)$ and assign 0 gives the solution $\bar{\lambda}=\frac{\varepsilon|D|^{2}\sigma^{2}}{M\tau^{2}}-\frac{1}{2}\geq 0$, and $h^{\prime\prime}(\bar{\lambda})>0$, implies $\lambda^{*}=\bar{\lambda}$. Plug into (E) gives:
@@
Theorem 6.2 is proved. ∎
Appendix F Experiments
F.1 Details of Datasets
@@
Synthetic. The synthetic data is generated using the same method in Li et al. [2020]. We briefly describe the generating steps here. For each device k, $y_{k}$ is computed from a softmax function $y_{k}=\text{argmax}(\text{softmax}(W_{k}x_{k}+b_{k}))$. $W_{k}$ and $b_{k}$ are drawn from the same Gaussian distribution with mean $u_{k}$ and variance $1$, where $u_{k}\in N(0;\beta)$. $x_{k}\in N(v_{k};\Sigma)$. $v_{k}$ is drawn from a Gaussian distribution with mean $B_{k}\in\mathcal{N}(0,\gamma)$ and variance $1$. $\Sigma$ is diagonal with $\sum j,j=j^{-1.2}$. In such a setting, $\beta$ controls how many local models differ from each other and $\gamma$ controls how much local data at each device differs from that of other devices.
In our experiment, we take $k=30$, $x\in\mathcal{R}^{20}$, $W\in\mathcal{R}^{10*20}$, $b\in\mathcal{R}^{10}$. We generate 4 datasets in total. They’re Syn(iid) Syn(0,0) with $\beta=0$ and $\gamma=0$, Syn(0.5,0.5) with $\beta=0.5$ and $\gamma=0.5$ and Syn(1,1) with $\beta=1$ and $\gamma=1$. In the output perturbation experiments, we set $\sigma$ to $1.0$ for the baseline methods and to $0.8$ for the Upcycled baselines, ensuring that the privacy budget $\epsilon$ for the baselines is always greater than that for the Upcycled baselines (e.g., the baseline $\bar{\epsilon}=0.773$ and the Upcycled baseline $\bar{\epsilon}=0.683$ for Syn(iid)). In the objective perturbation experiments, we set $\alpha$ to $10$ for the baselines and $20$ for the Upcycled baselines to achieve similar levels of information leakage, while still maintaining that $\epsilon$ for the Upcycled baselines is less than for the others. This constraint is maintained across all experiments.
FEMNIST: Similar with Li et al. [2020], we subsample 10 lower case characters (‘a’-‘j’) from EMNIST Cohen et al. [2017] and distribute 5 classes to each device. There are 50 devices in total. The input is 28x28 image. In the privacy experiments, $\sigma$ is set to $0.27$ for the baselines and $0.2$ for the Upcycled baselines, and $\alpha$ is set to $100$ for the baselines and $200$ for the Upcycled baselines, respectively.
Sent140: A text sentiment analysis task on tweets Go et al. [2009]. The input is a sequence of length 25 and the output is the probabilities of 2 classes. Here, $\sigma$ is set to $0.27$ for the baselines and $0.2$ for the Upcycled baselines, and $\alpha$ is set to $15$ and $30$ respectively.